OpenStudy (anonymous):

I have to figure out how many ounces of pesticide would be used in 3 gallons of water if the original parameters are one gallons of pesticide per for gallons of water. This is actually not a question for school but for work and I have to make sure I get the mix ratio right or burn up the plants. Thanks!

OpenStudy (anonymous):

one gallon of pesticide per 40 gallons of water..typo

OpenStudy (anonymous):

Burn Up The Plants?

OpenStudy (mathstudent55):

We are dealing with fluid ounces here. The conversion factor is 128 fluid ounces = 1 gallon.

OpenStudy (anonymous):

right. i know there are 128 fluid ounces in a gallon. I am just not sure how to set up the equation to make it work. I have been out of school for a couple years now. I need to start using my math again. I am forgetting lots of important stuff.

OpenStudy (mathstudent55):

The original ratio is: \( \dfrac{1 ~gal ~pesticide}{40 ~gal ~water} = \dfrac{128 ~fl ~oz~pesticide}{40 ~gal ~water} \)

OpenStudy (anonymous):

3.2 ounces per gallon of water?

OpenStudy (anonymous):

or 9.6 ounces per three gallons?

OpenStudy (mathstudent55):

Now you set the right fraction as a ratio in a proportion with the desired amount of water: \(\dfrac{128 ~fl ~oz~pesticide}{40 ~gal ~water} = \dfrac{x }{3 ~gal ~water} \)

OpenStudy (mathstudent55):

Correct.

OpenStudy (mathstudent55):

\( x = \dfrac{3~gal~water \times 128 ~fl ~oz~pesticide}{40 ~gal ~water} = 9.6 ~fl ~oz~pesticide\)

OpenStudy (anonymous):

thank you. I will write this down in my notepad for future reference. you saved the day!

OpenStudy (mathstudent55):

You are welcome. Good luck with the plants.